# Local order metrics for many-particle systems across length scales

### v1

@@ IV.2 Application of Number Variance-Based Order Metrics for Models in $d=1,2,3$
Figure 7 shows the scaled number variance $\sigma_{N}^{2}(R)/v_{1}(R)$ as a function of the scaled window radius $R\rho^{1/3}$ for each of the 3D models described in Sec. III. As above, the general qualitative characteristics of the curves follow from the discussion for $d=1$ and 2. Here, we note that across dimensions, the hard-core hyperuniform and nonhyperuniform system curves are clustered more tightly across the length scales considered is Fig. 7 due to the crystalline point patterns becoming relatively more disordered as $d$ increases, consistent with previous findings and the decorrelation principle [51, 89]. Similarly, HIP and HC-AHU have larger number density fluctuations at large $R$ for $d=3$ than for $d=2$.
Figure 8 shows $\Sigma_{N}(R)$ as a function of $R\rho^{1/3}$ for each of systems considered in Fig. 7. As above, qualitative statements regarding the relative rankings of different types of point patterns follow from the discussion of $\Sigma_{N}(R)$ for $d=1$ and 2. The key differences from $d=1$ and 2 include that, at small length scales, the FCC lattice is more ordered than the BCC lattice, which runs counter to the large-$R$ asymptotic result of BCC being more ordered (cf. the $B$ values of the FCC and BCC lattices in Table LABEL:tab:AB). Additionally, for $d=3$ HC-AHU is more ordered than URL, which is not the case in $d=1,2$, due to the range of its hard-core interactions increasing. At intermediate $R$, the high-$\chi$ stealthy hyperuniform point patterns are more ordered than $\mathbb{Z}^{3}$, consistent with their expected asymptotically large $R$ behavior given their respective $B$ values. This demonstrates that ordered hyperuniform point patterns are not necessarily the most ordered possible systems at a particular length scale, consistent with findings for two-phase media [25]. The HC-AHU system also now appears more disordered than the URL. In the neighborhood of intermediate values of $R$, the FCC and BCC curves intersect several times meaning one must go to much larger $R$ to match the asymptotic result of the BCC lattice minimizing $B$ in 3D (cf. Table LABEL:tab:AB) [88]. Moreover, the inset shows the gap between the $\mathbb{Z}^{3}$ lattice and the more ordered BCC and FCC lattices at large $R$, consistent with the notion that the $\mathbb{Z}^{d}$ lattices become less optimal (lower densities, worse solutions to the covering and quantizer problems) as $d$ increases compared to the optimal lattices in a given space dimension [88].
@@ V Discussion and Conclusions
In this work, we have devised order metrics to characterize the degree of order/disorder of many-body systems across all length scales via local number variance $\sigma_{N}^{2}(R)$. We found that the scaled number variance $\sigma_{N}^{2}(R)/v_{1}(R)$ and an integral measure derived from it $\Sigma_{N}(R_{i},R_{j})$ were able to sensitively characterize the degree of order/disorder across all length scales of 41 different point patterns that span antihyperuniform, standard nonhyperuniform, disordered hyperuniform, and ordered hyperuniform kinds with varying degrees of short-, intermediate-, and long-ranged order. Specifically, we found the degree of short-scale order is related directly to the degree of (anti)correlations in the small-$r$ regime of the $h(r)$ for a particular model; systems whose particles have exclusions regions ($h(r)=-1$ for $0<r<r^{\prime}$) have the greatest degree of small-scale order, while systems with $h(r)=+\infty$ as $r\rightarrow 0$ have the greatest degree of short-scale disorder. At intermediate length scales, we found that nonhyperuniform systems with hard-core interactions began to appear more disordered relative to the hyperuniform ones and antihyperuniform systems began to appear more disordered relative to nonhyperuniform systems. Finally, at large length scales, the aforementioned changes at intermediate length scales become more exaggerated. Across space dimensions, we find that a particular model becomes more disordered, consistent with the decorrelation principle [51]. Using $\Sigma_{N}(R)$, we showed that the ranking of the degree of order/disorder across models changes as a function of $R$, reaffirming that it is important to assess order/disorder with respect to a specific length scale [25, 26].
@@
The differences in length scale at which the nearly hyperuniform systems (e.g., RVL, equilibrium hard rods/disks/spheres) become markedly more disordered than hard-core hyperuniform systems are closely related to the ratio $B/A$, which quantifies the length scale at which a nearly hyperuniform system transitions from hyperuniform to nonhyperuniform $\sigma_{N}^{2}(R)$ scaling. The RVL is a particularly illuminating example because the defect concentration $p$ can be made small, resulting in large values of $B/A$. Thus, for sufficiently small $p$, RVLs can have hyperuniform $\sigma_{N}^{2}(R)$ scaling up to several orders of magnitude in $R\rho^{1/d}$, while having nonhyperuniform scaling on larger length scales. These systems have large and well-defined hyperuniform and nonhyperuniform scaling regimes, and thus, further exemplify the importance of assessing the degree of order/disorder with respect to a particular length scale [25, 26]. Our order metrics are able to sensitively detect the changes in order/disorder across length scales for the nearly hyperuniform systems examined herein, and we expect them to do so successfully for all other nearly hyperuniform point patterns for which $B/A$ is large (see, e.g., Ref. 46 for specific examples). In addition, at large length scales, we found that the rank-ordering of the nonhyperuniform and hyperuniform systems match the rankings of their $B$ and $A$ values, respectively.
@@
We note here that, while the number-variance based scalar order metrics can be applied to any many-particle system in $\mathbb{R}^{d}$, this is not the case of many other metrics including $s^{(2)}$ and $\tau$, given by Eqs. (20) and (21), respectively. In particular, $s^{(2)}$ and $\tau$ diverge when applied to systems with long-range order, such as lattices, vacancy-riddled lattices, or quasicrystals. This weakness can be addressed for the $\tau$ metric by using $\tau(K)$ [given by Eq. (22)], however, using $\tau(K)$ still requires one to incorporate a range of length scales $[2\pi/K,\infty)$ as opposed to one specific length scale. Moreover, they do not consistently treat antihyperuniform many-particle systems. For example, the HIP causes these metrics to diverge (i.e., indicating perfect order) because its $g_{2}(0)$ diverges, while the hard-core antihyperuniform system has $g_{2}(0)=0$ and is considered by $s^{(2)}$ and $\tau$ to have relatively little order. Moreover, all of the metrics discussed in Sec. IIC assess the degree of order/disorder of a many-particle system globally (i.e., not at some particular length scale) via some weighted integration of pair statistics, meaning one cannot determine at what specific length scales the system is ordered/disordered.
@@
Recalling the four properties of a “good” order metric from Sec. IIC, it is evident that $\tau$ and $s^{(2)}$ do not satisfy property 1 because of their inconsistent treatment of antihyperuniform systems. In addition, all existing metrics discussed in Sec. IIC are global assessments of order/disorder, and therefore do not satisfy property 3. In the present paper, we have shown that the number variance metrics satisfy all four properties of a good order metric set by Kansal et al. [5] (cf. Sec. IIC). Specifically, these metrics can rank the degree of order/disorder of systems with a broad spectrum of short-, intermediate-, and large-$R$ correlations consistently with physical intuition (properties 1 and 2) and, due to their explicit dependence on a length scale (property 3), are sensitive to local configurational patterns and the distribution thereof (property 4).
@@
To develop more sensitive order metrics, it would be valuable in future work to incorporate higher-order moments of the number density, such as those formulated in Ref. 43. The number variance metrics described in the present paper encode $\rho$ and $g_{2}(r)$ while higher-order moments as order metrics would encode not only $\rho$ and $g_{2}(r)$, but also $g_{3}(r)$, $g_{4}(r)$, etc. depending on the moment order. Thus, such metrics would additionally be sensitive to the relative distances between triplets, quartets, etc., of particles, allowing for a more comprehensive characterization of order/disorder for particle configurations. This would allow us to classify the order/disorder of systems with identical pair statistics but different three-body and higher-order correlations including, e.g., the pairs of equilibrium and nonequilibrium configurations with identical pair statistics from Ref. [90] to determine how the degree of order across length scales is affected by how the point pattern is generated.
@@ Appendix A Direct Sampling of the Number Variance
For systems that do not have analytical expressions for their pair statistics, we must directly sample the number density fluctuations in a point pattern to determine $\sigma_{N}^{2}(R)$, depicted schematically in Fig. 1. In this work, we follow the two-step procedure used in Ref. 46. First, observation windows are randomly (i.e., Poission) distributed throughout the point pattern. Then, we compute the number of points within each observation window using periodic boundary conditions. To reduce the computational load, the same set of window centers are used for all radii considered in a given point pattern. The number of windows is chosen such that the volume fraction of their union do not excess half the volume of the point pattern $V$, i.e., $0.51-\textrm{exp}[-N_{window}v_{1}(R_{max})/V]<$ 0.5, where $R_{max}$ is the largest window radius considered. This criterion follows from the exact formula for the volume fraction of Poisson-distributed overlapping spheres [83].

### v2

@@ IV.2 Application of Number Variance-Based Order Metrics for Models in $d=1,2,3$
Figure 7 shows the scaled number variance $\sigma_{N}^{2}(R)/v_{1}(R)$ as a function of the scaled window radius $R\rho^{1/3}$ for each of the 3D models described in Sec. III. As above, the general qualitative characteristics of the curves follow from the discussion for $d=1$ and 2. Here, we note that, across dimensions, the hard-core hyperuniform and nonhyperuniform system curves are clustered more tightly across the length scales considered in Fig. 7 due to the crystalline point patterns becoming relatively more disordered as $d$ increases, consistent with previous findings and the decorrelation principle [51, 89]. Similarly, HIP and HC-AHU have larger number density fluctuations at large $R$ for $d=3$ than for $d=2$.
Figure 8 shows $\Sigma_{N}(R)$ as a function of $R\rho^{1/3}$ for each of the systems considered in Fig. 7. As above, qualitative statements regarding the relative rankings of different types of point patterns follow from the discussion of $\Sigma_{N}(R)$ for $d=1$ and 2. The key differences from $d=1$ and 2 include that, at small length scales, the fcc lattice is more ordered than the bcc lattice, which runs counter to the large-$R$ asymptotic result of bcc being more ordered (cf. the $B$ values of the fcc and bcc lattices in Table LABEL:tab:AB). Additionally, for $d=3$ HC-AHU is more ordered than URL, which is not the case in $d=1,2$, due to the range of its hard-core interactions increasing. At intermediate $R$, the high-$\chi$ stealthy hyperuniform point patterns are more ordered than $\mathbb{Z}^{3}$, consistent with their expected asymptotically large $R$ behavior given their respective $B$ values. This demonstrates that ordered hyperuniform point patterns are not necessarily the most ordered possible systems at a particular length scale, consistent with findings for two-phase media [25]. The HC-AHU system also now appears more disordered than the URL. In the neighborhood of intermediate values of $R$, the fcc and bcc curves intersect several times meaning one must go to much larger $R$ to match the asymptotic result of the bcc lattice minimizing $B$ in 3D (cf. Table LABEL:tab:AB) [88]. Moreover, the inset shows the gap between the $\mathbb{Z}^{3}$ lattice and the more ordered bcc and fcc lattices at large $R$, consistent with the notion that the $\mathbb{Z}^{d}$ lattices become less optimal (lower densities, worse solutions to the covering and quantizer problems) as $d$ increases compared to the optimal lattices in a given space dimension [88].
@@ V Discussion and Conclusions
In this work, we have devised order metrics to characterize the degree of order/disorder of many-body systems across all length scales via the local number variance $\sigma_{N}^{2}(R)$. We found that the scaled number variance $\sigma_{N}^{2}(R)/v_{1}(R)$ and an integral measure derived from it $\Sigma_{N}(R_{i},R_{j})$ were able to sensitively characterize the degree of order/disorder across all length scales of 41 different point patterns that span antihyperuniform, standard nonhyperuniform, disordered hyperuniform, and ordered hyperuniform kinds with varying degrees of short-, intermediate-, and long-ranged order. Specifically, we found the degree of short-scale order is related directly to the degree of (anti)correlations in the small-$r$ regime of the $h(r)$ for a particular model; systems whose particles have exclusions regions [$h(r)=-1$ for $0<r<r^{\prime}$] have the greatest degree of small-scale order, while systems with $h(r)=+\infty$ as $r\rightarrow 0$ have the greatest degree of short-scale disorder. At intermediate length scales, we found that nonhyperuniform systems with hard-core interactions began to appear more disordered relative to the hyperuniform ones and antihyperuniform systems began to appear more disordered relative to nonhyperuniform systems. Finally, at large length scales, the aforementioned changes at intermediate length scales become more exaggerated. Across space dimensions, we find that a particular model becomes more disordered, consistent with the decorrelation principle [51]. Using $\Sigma_{N}(R)$, we showed that the ranking of the degree of order/disorder across models changes as a function of $R$, reaffirming that it is important to assess order/disorder with respect to a specific length scale [25, 26].
@@
The differences in length scale at which the nearly hyperuniform systems (e.g., RVL, equilibrium hard rods/disks/spheres) become markedly more disordered than hard-core hyperuniform systems are closely related to the ratio $B/A$, which quantifies the length scale at which a nearly hyperuniform system transitions from hyperuniform to nonhyperuniform $\sigma_{N}^{2}(R)$ scaling. The RVL is a particularly illuminating example because the defect concentration $p$ can be made small, resulting in large values of $B/A$. Thus, for sufficiently small $p$, RVLs can have hyperuniform $\sigma_{N}^{2}(R)$ scaling up to several orders of magnitude in $R\rho^{1/d}$, while having nonhyperuniform scaling on larger length scales. These systems have large and well-defined hyperuniform and nonhyperuniform scaling regimes and thus further exemplify the importance of assessing the degree of order/disorder with respect to a particular length scale [25, 26]. Our order metrics are able to sensitively detect the changes in order/disorder across length scales for the nearly hyperuniform systems examined herein and we expect them to do so successfully for all other nearly hyperuniform point patterns for which $B/A$ is large (see, e.g., Ref. 46 for specific examples). In addition, at large length scales, we found that the rank ordering of the nonhyperuniform and hyperuniform systems match the rankings of their $B$ and $A$ values, respectively.
@@
We note here that, while the number-variance based scalar order metrics can be applied to any many-particle system in $\mathbb{R}^{d}$, this is not the case of many other metrics including $s^{(2)}$ and $\tau$, given by Eqs. (20) and (21), respectively. In particular, $s^{(2)}$ and $\tau$ diverge when applied to systems with long-range order, such as lattices, vacancy-riddled lattices, or quasicrystals. This weakness can be addressed for the $\tau$ metric by using $\tau(K)$ [given by Eq. (22)]; however, using $\tau(K)$ still requires one to incorporate a range of length scales $[2\pi/K,\infty)$ as opposed to one specific length scale. Moreover, they do not consistently treat antihyperuniform many-particle systems. For example, the HIP causes these metrics to diverge (i.e., indicating perfect order) because its $g_{2}(0)$ diverges, while the hard-core antihyperuniform system has $g_{2}(0)=0$ and is considered by $s^{(2)}$ and $\tau$ to have relatively little order. Moreover, all of the metrics discussed in Sec. IIC assess the degree of order/disorder of a many-particle system globally (i.e., not at some particular length scale) via some weighted integration of pair statistics, meaning one cannot determine at what specific length scales the system is ordered/disordered.
@@
Recalling the four properties of a “good” order metric from Sec. IIC, it is evident that $\tau$ and $s^{(2)}$ do not satisfy property (1) because of their inconsistent treatment of antihyperuniform systems. In addition, all existing metrics discussed in Sec. IIC are global assessments of order/disorder and therefore do not satisfy property (3). In the present paper, we have shown that the number variance metrics satisfy all four properties of a good order metric set by Kansal et al. [5] (cf. Sec. II C). Specifically, these metrics can rank the degree of order/disorder of systems with a broad spectrum of short-, intermediate-, and large-$R$ correlations consistently with physical intuition [properties (1) and (2)] and, due to their explicit dependence on a length scale [property (3)], are sensitive to local configurational patterns and the distribution thereof [property (4)].
@@
To develop more sensitive order metrics, it would be valuable in future work to incorporate higher-order moments of the number of points within a spherical window of radius $R$, such as those formulated in Ref. 43. The number variance metrics described in the present paper encode $\rho$ and $g_{2}(r)$, while higher-order moments as order metrics would encode not only $\rho$ and $g_{2}(r)$, but also $g_{3}(r)$, $g_{4}(r)$, etc., depending on the moment order. Thus such metrics would additionally be sensitive to the relative distances between triplets, quartets, etc., of particles, allowing for a more comprehensive characterization of order/disorder for particle configurations. This would allow us to classify the order/disorder of systems with identical pair statistics but different three-body and higher-order correlations including, e.g., the pairs of equilibrium and nonequilibrium configurations with identical pair statistics from Ref. [90] to determine how the degree of order across length scales is affected by how the point pattern is generated.
@@ Appendix A Direct Sampling of the Number Variance
For systems that do not have analytical expressions for their pair statistics, we must directly sample the number density fluctuations in a point pattern to determine $\sigma_{N}^{2}(R)$, depicted schematically in Fig. 1. In this work, we follow the two-step procedure used in Ref. 46. First, observation windows are randomly (i.e., Poission) distributed throughout the point pattern. Then, we compute the number of points within each observation window using periodic boundary conditions. To reduce the computational load, the same set of window centers are used for all radii considered in a given point pattern. The number of windows is chosen such that the volume fraction of their union doed not exceed half the volume of the point pattern $V$, i.e., $0.51-\textrm{exp}[-N_{window}v_{1}(R_{max})/V]<$ 0.5, where $R_{max}$ is the largest window radius considered. This criterion follows from the exact formula for the volume fraction of Poisson-distributed overlapping spheres [83].